find the sum of the first n terms of each arithmetic series
n=18; a1=40, d=-7

Answer by checkley77(12844)

First we need to find the last number in the 18 series.
L=a+(n-1)D
L=40+(18-1)-7
L=40+17*-7
L=40-119
L=-49 this is the 18 th. number.
Sum=n/2(a+L)
Sum=18/2(40-79)
Sum=9*39
Sum=-351 ans.